A union worker made $500 per week. His union sought a one-year contract and went on strike. Once the new contact was approved, it provided for a 4% raise. Assume that the worker was not paid during the strike. Given his raise in salary, how many weeks could he strike and still make at least as much for the next 52 weeks as he would have made without a strike?

Answer by ReadingBoosters(3246)

Pre-strike
500 * 52 = 26,000 per year
With $500 a week for 52 weeks, he earns a salary of $26,000 per year.
...
Post-strike
4% (26,000) = .04(26,000) = 1,040 raise
4% of his salary, $26,000, yields a $1,040 raise.
26,000 + 1,040 = 27,040 per year
His raise aded to his salary yields a new salary of $27,040
27,040/52 = $520/wk
Taking his new salary over the next 52 weeks yields a weekly pay of $520 per week.
1040/520 = 2 weeks
The amount of raise divided by weekly pay suggests how long before the raise is used.
He could strick for 2 weeks.
...
Proof
If on strike for 2 weeks, salary has been reduced from 52 weeks to 50 weeks at the rate of $520 per week.
520(50) = 26,000, which is the same as the pre-strike salary.
.......................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.